40. Calculation of drug dose concentration Example Vials of Penicillin are available in 5 milliliter suspensions of 10 mg/ml concentration. You are asked to inject 1 mg of penicillin to a patient intramuscularly. What volume of the suspension in the vial you would give to the patient?
Answer #1

ANSWER

40. Drug dosage concentration.

               Formula

   Volume of suspension give to the patient = Amount desired(D) / Amount on hand(H) * Quantity(Q)

         Amount desired (D) = 1mg.

        Amount in hand (H) = 10 mg.

              Quantity = 1 ml.

   Volume of suspension required for the patient = D / H * Q

                                                                             = 1 / 10 *1

                                                                               = 0.1 * 1

                                                                              = 0.1 ml.